# Scalable LLM Math Reasoning Acceleration with Low-rank Distillation

### v2

@@ Appendix C LoRA Performance
GRIFFIN 7.13 5.42 56.05 14.11
Layer-wise LoRA 8.11 5.60 56.00 14.65
E2E LoRA 20.47 10.10 56.10 14.41
Layer-wise Caprese 13.72 6.62 56.07 13.40
E2E Caprese 21.00 8.44 56.55 13.88
@@
GRIFFIN 28.96 10.98 64.52 12.52
Layer-wise LoRA 26.46 11.28 63.88 12.55
E2E LoRA 42.91 16.50 64.67 12.63
Layer-wise Caprese 40.18 13.70 64.33 11.60
E2E Caprese 44.66 16.96 64.83 12.35
Gemma 2 2B Instruct 51.02 16.06 63.77 10.96
GRIFFIN 33.74 11.32 63.28 11.07
Layer-wise LoRA 39.58 14.66 62.98 10.54
E2E LoRA 44.66 16.71 63.03 10.81
Layer-wise Caprese 42.53 12.32 63.77 10.75
E2E Caprese 48.14 13.70 63.37 11.05
Table 5: 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA and QASPER) with LoRA. $F$ sparsity is set to 50%, $r=256$, and LoRA ranks are set to match the number of parameters that Caprese adds.
Appendix D GRIFFIN Response Lengths
@@
Appendix E Training Hyperparameters
@@
Table 6: Caprese layer-wise and E2E training hyperparameters.

### v3

@@ Appendix D Caprese with Quantization
QLoRA 12.05 6.56 48.70 12.37
Layer-wise Caprese 15.01 6.88 49.60 12.65
E2E Caprese 21.61 9.98 53.52 13.54
@@
QLoRA 45.41 11.98 62.22 12.87
Layer-wise Caprese 45.56 11.78 64.25 12.41
E2E Caprese 48.82 14.14 64.05 12.26
Appendix E GRIFFIN Response Lengths
@@
Appendix F Rank of Learned Parameters
In Figure 8, we plot the relative singular values for each learned product $\bm{L}\bm{R}$ from (7) in Llama 3.2 3B Instruct. Although we set the inner rank to be 256, we see that some layers in Caprese(CATS) are still very low rank. In comparison, Caprese(GRIFFIN) layers are relatively high rank, likely to due to the fact that GRIFFIN performs highly structured FF sparsification. From this, we hypothesize that GRIFFIN may better utilize an increased Caprese inner dimension.
Figure 8: Relative singular values of learned Llama 3.2 3B Instruct Caprese layers. Blue lines are individual layers; red lines are averaged across layers.
Appendix G Training Hyperparameters
@@
Table 8: Caprese layer-wise and E2E training hyperparameters.